### v1

# Emotions Where Art Thou: Understanding and Characterizing the Emotional Latent Space of Large Language Models

Abstract
This work investigates how large language models (LLMs) internally represent emotion by analyzing the geometry of their hidden-state space. The paper identifies a low-dimensional emotional manifold and shows that emotional representations are directionally encoded, distributed across layers, and aligned with interpretable dimensions. These structures are stable across depth and generalize to eight real-world emotion datasets spanning five languages. Cross-domain alignment yields low error and strong linear probe performance, indicating a universal emotional subspace. Within this space, internal emotion perception can be steered while preserving semantics using a learned intervention module, with especially strong control for basic emotions across languages. These findings reveal a consistent and manipulable affective geometry in LLMs and offer insight into how they internalize and process emotion.
1 Introduction
Large Language Models (LLMs) have become central tools for interacting with, analyzing, and generating human language. Their widespread deployment across domains has led to increasing interest in how they handle not just syntactic or semantic meaning, but also affective tone. Emotion is a fundamental part of language, shaping persuasion, social signaling, and narrative context. As such, understanding how LLMs process emotional content is essential for both interpretability and safe deployment.
The literature on affect in NLP has focused on sentiment analysis, a task where models classify inputs into discrete emotional or affective categories Prabowo & Thelwall (2009); Medhat et al. (2014); Wadawadagi & Pagi (2020); Wankhade et al. (2022). While this demonstrates that LLMs can identify emotions, it offers little insight into how emotional meaning is represented internally. Classification accuracy is not equivalent to interpretability.
Other works have taken a behavioral view, exploring the emotional “intelligence” of LLMs. These include prompting models with hypothetical emotional scenarios and evaluating their responses Wang et al. (2023), or probing how well they align with human judgments in affective tone Huang et al. (2024); Zhao et al. (2023). Though these studies suggest some degree of affective sensitivity, they focus on testing outputs rather than investigating internal mechanisms of the LLM.
Recent work has also examined emotion manipulation and decoding. For instance, models have been used to map text to dimensional emotion ratings like valence-arousal-dominance (VAD) Shah et al. (2023); Broekens et al. (2023), or to generate emotionally inflected language on demand Reichman et al. (2025). LLMs have also been shown to be more likely to comply with emotionally framed requests Vinay et al. (2024). These studies also treat emotion primarily as a label or generation condition—not a latent internal representation.
While there has been some work examining how LLMs respond to or generate emotional language, the structure of emotional representations within their hidden states remains relatively underexplored. Most prior approaches focus on output behavior or classification accuracy, with comparatively few efforts aimed at interpreting the internal geometry of emotion encoding. To advance the understanding of how emotions are represented in LLMs and how they influence LLM responses, this paper investigates the internal mechanisms of the LLM. Emotions are analyzed within LLM hidden states across layers, datasets, and languages.
Our contributions are as follows: (1) We extract a low-dimensional emotional subspace of the LLM and show that it captures interpretable, directionally encoded affective structure across LLM layers. (2) We demonstrate that this space generalizes across eight emotion datasets spanning five languages, with low alignment distortion and high cross-domain probe accuracy. (3) We introduce a learned steering module that manipulates internal emotional perception while preserving semantic content, with especially strong control over basic emotions. We find that emotional encoding is directional, distributed, and remarkably consistent across varied textual modalities. We also investigate the model’s internal “psychology”: how emotions are separated, aligned, and—critically—how they can be steered via targeted interventions.
2 Related Works
Models of Emotions. Psychological models of emotion are commonly categorized as either discrete or continuous. Discrete theories posit that emotions are fundamentally distinct categories—such as the six “basic” emotions proposed by Ekman et al. (1999): anger, surprise, disgust, enjoyment, fear, and sadness. Other taxonomies expand this set, including more nuanced affective states Plutchik (1991).
In contrast, continuous models view emotions as points in a low-dimensional latent space. A widely used formulation is the valence-arousal-dominance (VAD) model Mehrabian (1996), where valence encodes hedonic tone, arousal measures intensity, and dominance reflects control or agency. Variants of this framework reduce or alter the axes (e.g., Russell’s 2D circumplex Russell (1980)).
These representations offer an interpretive lens for analyzing learned emotion structure in LLMs: If models implicitly encode emotions in a geometric space, we may expect that certain latent directions align with these classic dimensions. Our work explores whether such a structure emerges naturally in the hidden-state geometry of LLMs trained without explicit emotional supervision.
Neuroscientific models of emotion offer a parallel debate. Localist theories posit that discrete emotions correspond to specific, anatomically distinct brain regions, while constructionist theories argue that emotions emerge from distributed, domain-general processes Lindquist et al. (2012); Vytal & Hamann (2010); Celeghin et al. (2017). Our results, particularly from ML-AURA (Section 5), support a constructionist-style interpretation in LLMs: emotional content is not localized to a small subset of units but is instead widely distributed across neurons and layers, with high separability emerging from overlapping, multi-purpose components.
Emotions in Latent Space. Recent work has investigated how LLMs interact with emotional text, often focusing on behavior or output-level mappings. For example, ChatGPT has shown the ability to map emotions to Valence-Arousal-Dominance (VAD) values Broekens et al. (2023); Yongsatianchot et al. (2023), suggesting that emotion-relevant dimensions are accessible to the model. However, such studies do not analyze the internal structure or geometry of these latent representations.
Some prior work explicitly trains models to embed emotions into structured spaces, using classification objectives or external supervision. For instance, Dathathri et al. (2019) and Buechel et al. (2020) train models to map between emotion spaces. Similarly, Wang & Zong (2021) learns an emotion space from labeled data, shows clustering by valence, and demonstrates transferability across datasets. However, in all of these works, the emotion space is imposed or supervised, not emergent.
A growing line of work probes how pretrained models encode emotion. Hollinsworth et al. (2024) show that valence is linearly embedded in contextual states, while Zhang et al. (2023) find arousal and dominance less separable, though their analysis depends on encoder-only models and fixed affective lexicons. In contrast, we study decoder-only LLMs, seeking to recover emergent emotional structure directly from hidden-state geometry rather than imposing a psychological model.
Other studies have shown that LLMs exhibit strong zero-shot emotion classification performance across languages Bianchi et al. (2022), though subsequent work notes that language-specific tuning is sometimes necessary for culturally grounded affect De Bruyne et al. (2022). These findings suggest that emotion representations are at least partially transferable across linguistic domains—a hypothesis we test more directly through geometric alignment and projection-based analysis in Section 4.
3 Methods
To understand how emotions are represented in LLMs, a variety of tools were used. This section outlines those methods and their theoretical grounding. Empirical findings from these analyses are presented in Sections 4 and 5.
ML-AURA. ML-AURA quantifies how selectively a neuron responds to a specific concept by framing each neuron as a threshold-based detector Suau et al. (2024). For a labeled dataset $D$, each neuron’s output is summarized per example using the maximum activation across tokens. These scalar responses are then ranked and evaluated using the area under the precision-recall curve, comparing neuron output against the presence or absence of the target concept. Neurons with high AUC-PR are designated as “experts” for that concept.
In our adaptation, the concepts are emotion categories. We apply ML-AURA in a one-vs-all setup for each emotion, scoring each neuron by how well it distinguishes a target emotion from all others.
Centered-SVD. We build on prior work showing that LLM hidden states lie on low-dimensional manifolds where semantic and syntactic properties are linearly recoverable Aghajanyan et al. (2020); Hu et al. (2022); Lizzo & Heck (2025). To isolate emotion-relevant subspaces, we apply singular value decomposition (SVD) to hidden state activations.
We use the dataset from Reichman et al. (2025), which pairs neutral sentences with synthetic rewrites expressing target emotions. Sentence-level hidden states are obtained by mean-pooling token activations. The resulting activation matrix is centered and decomposed with SVD, yielding principal directions in representation space. Since emotional content is the primary structured variation across inputs, we hypothesize that the leading components capture dominant emotional axes. We test this through alignment, probing, and causal manipulation.
Space Alignment. Prior work has shown that latent spaces arising from related tasks often exhibit similar internal geometry, with relationships between them approximately rigid or linear up to rescaling and rotation Moschella et al. (2023). While some approaches lift these spaces into anchor-relative representations to handle isometric variance, recent work demonstrates that direct alignment via linear or rigid transformations is often sufficient and easier to apply in practice Lähner & Moeller (2024). Following this approach, we use linear regression to align the emotional subspace derived from synthetic data with that derived from human-authored emotion classification datasets. This alignment allows us to test whether the structure found in the synthetic manifold reflects transferable emotional encodings or artifacts specific to the synthetic generation process.
4 Emotion Universality
Using the tools presented in Section 3, we provide evidence that emotional representations in LLMs are structurally universal. We show that emotions are encoded in similar geometric subspaces across datasets, languages, and writing styles. This and all subsequent sections focus on LLaMA 3.1; analogous results for Olmov2 and Ministral are provided in the appendix.
4.1 Datasets
We evaluate the universality of emotional representations in LLMs using eight diverse datasets, each offering explicit categorical emotion labels; datasets restricted to polarity or star ratings were excluded as too coarse. The collection spans languages, modalities, and styles: Go-Emotions contains English Reddit comments Demszky et al. (2020), CARER covers English tweets Saravia et al. (2018), SemEval-2007 Task 14 focuses on English news headlines Strapparava & Mihalcea (2007), EmoEvent includes English and Spanish tweets Plaza-del-Arco et al. (2020), Emotions in Drama consists of German plays from the eighteenth and nineteenth centuries Dennerlein et al. (2023), Bhaav offers Hindi short stories Kumar et al. (2019), MultiEmotions-It provides Italian YouTube and Facebook comments Sprugnoli et al. (2020), and EmoTextToKids features French journalistic and encyclopedic texts written for children Étienne et al. (2024). Appendix A analyzes the structural and stylistic contents of each of these datasets.
The chosen languages are those for which high-quality emotion datasets exist and which are officially supported by LLaMA 3.1, as specified in its technical report.
4.2 Universality Analysis
The first step of the universality analysis was to collect the mean-pooled hidden-states of the model when text $x_{i}$ from dataset $D$ is input into it. Then the mean-pooled hidden state is either projected onto the space described in Section 3 or statistics are directly derived from comparing the mean-pooled hidden states.
Dataset Stress-1 $\downarrow$ Stress-2 $\downarrow$ Sammon $\downarrow$ Avg Dist $\downarrow$ $\ell_{2}$ $\downarrow$ $\sigma$ $\downarrow$ Probe Acc. $\uparrow$
Llama3.1-8B-Base
Go-Emotions 0.33 ± 0.33 0.13 ± 0.13 43%* 0.52 ± 0.52
CARER (Twitter) 0.38 ± 0.16 0.17 ± 0.15 0.18 ± 0.22 1.03 ± 0.24 1.11 ± 0.32 0.16 ± 0.33 0.35 ± 0.11
SemEval 0.34 ± 0.16 0.14 ± 0.15 0.14 ± 0.19 0.97 ± 0.23 1.03 ± 0.31 0.12 ± 0.26 0.46 ± 0.13
EmoEvent (EN) 0.36 ± 0.13 0.15 ± 0.13 0.15 ± 0.2 0.9 ± 0.19 0.97 ± 0.29 0.14 ± 0.26 0.55 ± 0.12
EmoEvent (ES) 0.38 ± 0.14 0.17 ± 0.14 0.16 ± 0.22 0.86 ± 0.2 0.93 ± 0.31 0.16 ± 0.29 0.5 ± 0.15
Bhaav (Hindi) 0.39 ± 0.15 0.17 ± 0.15 0.17 ± 0.19 0.86 ± 0.22 0.92 ± 0.31 0.16 ± 0.31 0.36 ± 0.1
German Drama 0.39 ± 0.17 0.18 ± 0.18 0.22 ± 0.47 1.06 ± 0.22 1.17 ± 0.48 0.48 ± 4.27 0.29 ± 0.1
MultiEmotions-It 0.38 ± 0.16 0.17 ± 0.15 0.19 ± 0.2 1.11 ± 0.24 1.19 ± 0.28 0.18 ± 0.48 0.46 ± 0.12
EmoTextToKids (FR) 0.41 ± 0.14 0.19 ± 0.15 0.19 ± 0.19 0.92 ± 0.24 0.99 ± 0.32 0.18 ± 0.33 0.39 ± 0.1
Average (Full-Space) 0.38 ± 0.15 0.17 ± 0.15 0.17 ± 0.25 0.97 ± 0.24 1.04 ± 0.35 0.2 ± 1.55 0.42 ± 0.14
Average (50D-Space) 0.58 ± 0.13 0.35 ± 0.16 0.33 ± 0.24 0.6 ± 0.21 0.68 ± 0.36 0.3 ± 1.66 0.39 ± 0.11
Llama3.1-8B-Instruct
Go-Emotions 0.41 ± 0.11 0.18 ± 0.17 71%* 0.42 ± 0.06
CARER (Twitter) 0.4 ± 0.11 0.17 ± 0.11 40%* 0.58 ± 0.07
SemEval 0.6 ± 0.05 0.36 ± 0.06 0.32 ± 0.07 0.52 ± 0.09 0.56 ± 0.2 0.15 ± 0.28 0.49 ± 0.09
EmoEvent (EN) 0.4 ± 0.1 0.17 ± 0.09 0.2 ± 0.22 1.04 ± 0.14 1.15 ± 0.28 0.2 ± 0.3 0.11 ± 0.01
EmoEvent (ES) 0.43 ± 0.1 0.19 ± 0.1 0.22 ± 0.23 1.0 ± 0.14 1.12 ± 0.3 0.24 ± 0.38 0.11 ± 0.01
Bhaav (Hindi) 0.43 ± 0.11 0.2 ± 0.11 0.23 ± 0.3 1.02 ± 0.17 1.14 ± 0.38 0.23 ± 0.26 0.55 ± 0.05
German Drama 0.51 ± 0.15 0.29 ± 0.29 71%* 0.43 ± 0.07
MultiEmotions-It 0.49 ± 0.12 0.25 ± 0.13 46%* 0.57 ± 0.07
EmoTextToKids (FR) 0.42 ± 0.1 0.18 ± 0.1 0.21 ± 0.22 1.02 ± 0.13 1.13 ± 0.3 0.22 ± 0.27 0.60 ± 0.08
Average (Full-Space) 0.46 ± 0.12 0.22 ± 0.12 0.24 ± 0.23 0.92 ± 0.24 1.02 ± 0.37 0.21 ± 0.3 0.43 ± 0.19
Average (50D-Space) 0.55 ± 0.11 0.31 ± 0.13 0.33 ± 0.23 0.83 ± 0.27 0.96 ± 0.42 0.33 ± 0.35 0.37 ± 0.16
Olmov2-Base
Go-Emotions 0.71 ± 0.36 0.63 ± 0.98 1.39 ± 5.32 1.65 ± 0.72 1.94 ± 1.08 0.34 ± 0.25 0.48 ± 0.06
CARER (Twitter) 0.66 ± 0.36 0.57 ± 1.03 1.22 ± 4.97 1.54 ± 0.67 1.8 ± 1.0 0.33 ± 0.25 0.61 ± 0.06
SemEval 0.78 ± 0.41 0.78 ± 0.95 1.35 ± 2.71 1.8 ± 0.63 2.02 ± 0.84 0.25 ± 0.18 0.6 ± 0.06
EmoEvent (EN) 0.56 ± 0.23 0.36 ± 0.42 0.75 ± 2.49 1.38 ± 0.5 1.61 ± 0.8 0.32 ± 0.24 0.11 ± 0.01
EmoEvent (ES) 0.55 ± 0.25 0.37 ± 0.5 0.76 ± 2.84 1.31 ± 0.59 1.53 ± 0.9 0.32 ± 0.22 0.11 ± 0.01
Bhaav (Hindi) 0.61 ± 0.26 0.44 ± 0.47 0.92 ± 2.86 1.26 ± 0.72 1.55 ± 1.09 0.44 ± 0.28 0.51 ± 0.05
German Drama 0.77 ± 0.58 0.93 ± 3.16 2.1 ± 12.92 1.42 ± 1.08 1.78 ± 1.53 0.52 ± 0.3 0.44 ± 0.05
MultiEmotions-It 0.74 ± 0.57 0.87 ± 2.74 1.88 ± 10.48 1.39 ± 1.04 1.71 ± 1.43 0.48 ± 0.28 0.58 ± 0.06
EmoTextToKids (FR) 0.6 ± 0.27 0.43 ± 0.59 0.87 ± 3.18 1.36 ± 0.47 1.62 ± 0.85 0.37 ± 0.33 0.59 ± 0.06
Average (Full-Space) 0.67 ± 0.4 0.6 ± 1.56 1.25 ± 6.43 1.46 ± 0.76 1.73 ± 1.1 0.37 ± 0.27 0.45 ± 0.2
Average (50D-Space) 0.72 ± 0.4 0.68 ± 1.61 1.56 ± 8.11 1.37 ± 0.9 1.78 ± 1.41 0.63 ± 0.42 0.34 ± 0.14
Olmov2-Instruct
Go-Emotions 0.55 ± 0.19 0.34 ± 0.31 63%* 0.49 ± 0.06
CARER (Twitter) 0.57 ± 0.21 0.37 ± 0.41 38%* 0.65 ± 0.08
SemEval 0.46 ± 0.21 0.25 ± 0.3 47%* 0.61 ± 0.09
EmoEvent (EN) 0.51 ± 0.18 0.3 ± 0.26 41%* 0.11 ± 0.01
EmoEvent (ES) 0.51 ± 0.2 0.3 ± 0.3 41%* 0.11 ± 0.01
Bhaav (Hindi) 0.61 ± 0.27 0.44 ± 0.6 78%* 0.53 ± 0.05
German Drama 0.66 ± 0.29 0.52 ± 0.72 61%* 0.43 ± 0.05
MultiEmotions-It 0.65 ± 0.31 0.52 ± 0.88 41%* 0.56 ± 0.05
EmoTextToKids (FR) 0.56 ± 0.23 0.37 ± 0.45 34%* 0.62 ± 0.07
Average (Full-Space) 0.56 ± 0.24 0.38 ± 0.52 49%* 0.46 ± 0.20
Average (50D-Space) 0.65 ± 0.25 0.48 ± 0.58 979.03 ± 3086.26 882.19 ± 3055.12 482470.52 ± 978732.71 2673197.92 ± 3939472.37 0.36 ± 0.15
Ministral
Go-Emotions 0.45 ± 0.17 0.23 ± 0.35 68%* 0.4 ± 0.05
CARER (Twitter) 0.47 ± 0.11 0.23 ± 0.11 70%* 0.49 ± 0.08
SemEval 0.38 ± 0.2 0.19 ± 0.51 0.24 ± 0.86 1.13 ± 0.19 1.22 ± 0.37 0.16 ± 0.49 0.55 ± 0.07
EmoEvent (EN) 0.41 ± 0.11 0.18 ± 0.1 0.21 ± 0.2 1.1 ± 0.14 1.2 ± 0.23 0.18 ± 0.15 0.11 ± 0.01
EmoEvent (ES) 0.48 ± 0.12 0.25 ± 0.12 0.32 ± 0.26 1.24 ± 0.14 1.37 ± 0.25 0.2 ± 0.16 0.11 ± 0.01
Bhaav (Hindi) 0.42 ± 0.11 0.19 ± 0.10 75%* 0.45 ± 0.05
German Drama 0.50 ± 0.25 0.32 ± 0.98 70%* 0.45 ± 0.07
MultiEmotions-It 0.49 ± 0.12 0.26 ± 0.14 0.31 ± 0.21 1.18 ± 0.11 1.31 ± 0.18 0.24 ± 0.26 0.66 ± 0.05
EmoTextToKids (FR) 0.43 ± 0.11 0.2 ± 0.11 0.24 ± 0.22 1.13 ± 0.12 1.24 ± 0.22 0.19 ± 0.16 0.57 ± 0.07
Average (Full-Space) 0.45 ± 0.16 0.23 ± 0.39 0.26 ± 0.44 1.16 ± 0.15 1.27 ± 0.27 0.19 ± 0.28 0.42 ± 0.19
Average (50D-Space) 0.52 ± 0.14 0.29 ± 0.4 0.32 ± 0.45 1.04 ± 0.18 1.19 ± 0.31 0.31 ± 0.3 0.36 ± 0.16
Table 1: Per-dataset distortion metrics and probe accuracy across three models. Lower distortion indicates greater geometric consistency. * in cells denotes very high stress/distortion. Instead of reporting the stress or distortion for that dataset, the percentage of layers that are highly distorted are reported.
Model Avg Cosine Similarity Avg MSE Avg Spectral Flatness Avg Frob Norm
Llama3.1-8B-Base 0.84 ± 0.12 1.79 ± 1.97 2.09 ± 0.39 7.70 ± 2.67
Llama3.1-8B-Instruct 0.93 ± 0.06 0.91 ± 1.07 2.29 ± 0.42 8.68 ± 6.36
Olmov2-Base 0.86 ± 0.15 1.87 ± 5.34 2.28 ± 0.40 8.02 ± 1.62
Olmov2-Instruct 0.90 ± 0.10 1.00 ± 2.17 2.26 ± 0.40 7.98 ± 1.31
Ministral 0.92 ± 0.04 1.44 ± 1.47 2.22 ± 0.44 7.60 ± 0.91
Table 2: Layer-wise average cosine similarity and linear alignment statistics by model.
Table 2 reports cosine similarity, regression error, and transformation statistics, each computed at every sublayer and then averaged across the full depth of the model. All models achieve a strong average centroid cosine similarity between emotions in real datasets and their synthetic counterparts (0.84–0.93 on average), showing that emotional categories are encoded in similar directions across domains and languages. Instruction-tuned models, especially LLaMA-3.1-Instruct and OLMo-v2-Instruct, show higher similarity and lower MSE than their base versions, indicating that tuning improves how closely real-emotion spaces align with the synthetic manifold. By contrast, spectral flatness and Frobenius norm values are similar across models and do not reliably separate base from instruct variants: LLaMA shows a slight increase with tuning, OLMo shows none or a decrease, and Ministral lies in the same range. This suggests that the improved alignment is not due to stronger or more isotropic transformations, but rather to better-aligned underlying representations.
With the alignment between emotional spaces established, the distortion scores for each space of each model are assessed next. Results are summarized in Table  1. While traditional usage defines Stress-1 below 0.2 as acceptable and below 0.1 as strong Kruskal (1964), these thresholds were developed for low-dimensional embeddings (e.g., 2D). In the setting described here—mapping between high-dimensional hidden states—no canonical thresholds exist, but consistently low scores indicate robust relational preservation. This is most clearly observed in LLaMA-3.1-8B-Base, which achieves the lowest stress values across datasets. LLaMA-3.1-8B-Instruct and Ministral also yield relatively low Stress-1 and Stress-2 scores, though somewhat higher than the LLaMA-3.1-Base model, suggesting that instruction-tuned variants preserve emotional relations slightly less consistently. Interestingly, OLMo-v2 shows the opposite trend: the base model achieves lower stress than the instruction-tuned variant, but both versions exhibit substantially higher stress values than LLaMA or Ministral. This pattern suggests that OLMo encodes a less unified emotional space than the other models. Projection into the 50D synthetic subspace increases stress across all models, consistent with expected compression effects. For models with already low stress values, the increases remain modest and the geometry remains coherent, whereas for models with elevated stress (e.g., OLMo), dimensionality reduction does not mitigate distortion.
To complement the stress scores, three high-dimensional embedding distortion metrics are reported: average distortion, $\ell_{2}$-distortion, and $\sigma$-distortion. Values near 1 for the first two indicate that pairwise distances are preserved up to global scaling, while low $\sigma$-distortion (ideal value =0=0) reflects stable proportionality in relative distances. These metrics highlight the contrast between base and instruction-tuned models. Both LLaMA-3.1-8B-Base and OLMo-v2-Base achieve distortion scores near ideal across most datasets (recall that no Ministral-Base model exists). For LLaMA-3.1-8B-Base, only two datasets deviate substantially from the ideal; excluding these outliers, its distortion metrics are more favorable than those of OLMo-v2-Base in both the full hidden-state space and the 50D synthetic subspace. See Appendix C for a contextualization of these scores.
The instruction-tuned models, however, diverge considerably from their base counterparts. LLaMA-3.1-8B-Instruct produces roughly twice as many outlier datasets as LLaMA-3.1-8B-Base, while Ministral exhibits five “outliers” out of the nine datasets tested. Even after excluding these outliers, the distortion metrics for both models remain less favorable than those of their base equivalents in both full-space and 50D-space analyses, though still reasonably close to the ideal range. Nonetheless, for LLaMA-3.1-8B-Instruct and Ministral, the emotional embedding space remains partially universal: emotions are still represented in a broadly consistent manner across multiple datasets and languages. However, OLMo-v2-Instruct performs markedly worse. The spaces formed by emotion activations across different datasets are essentially mutually unintelligible. Its average distortion scores are extremely high across the entire model, suggesting that each dataset encodes emotions in a dataset- and language-specific fashion that resists mapping into a shared space.
Rather than reporting average distortion per layer for datasets with elevated distortion, we instead report the percentage of layers exhibiting high distortion. In LLaMA-3.1-8B-Base, only Go-Emotions shows substantial distortion (43% of layers). The instruction-tuned variant shows broader distortion on Go-Emotions, though across datasets a majority of layers are affected only about half the time. OLMo-v2-Instruct shows majority distortion in 3 of 9 datasets, with the rest under 50% (some under 40%). Ministral is most fragile: when distortion appears, it usually spans most layers, though fewer datasets are affected. Together, these results indicate that even when a model exhibits distortion in parts of its architecture, large fractions of its layers still encode emotional spaces that remain universal across datasets, languages, and writing styles. These layers likely play a central role in maintaining consistent emotional representation within the model.
Linear probes trained on activations projected into the synthetic emotional latent space predict the emotions of human-written text at above-chance levels. These results indicate that emotional structure in LLMs is not only geometrically consistent but also linearly decodable across diverse domains, though probe performance varies with model family and the coherence of the underlying emotional space.
Comparing Table 2 with Table 1 underscores a central tension. The layer-averaged cosine similarities and regression errors suggest that all models align well with the synthetic emotional manifold, often with small differences across families. Yet the stress and distortion metrics reveal that, within the same models, relational structure can still be substantially warped—sometimes across large fractions of layers. This discrepancy reflects the fact that centroidal and regression-based measures capture global alignment, whereas stress and distortion expose finer-grained deviations in how relative distances between emotions are preserved. Thus, high apparent alignment at the aggregate level can coexist with local irregularities in the geometry of emotional spaces; however, linear probing shows that these spaces can still usefully and predictably predict emotion, even if there is some distortion between the synthetic and human-written spaces.
Figure 1: Results of ML-AURA by layer and emotion. Results are in terms of percent of neurons with an AUROC score above $0.9$.
5 Model Psychology
Having established the external consistency of emotional geometry across datasets, we now turn inward, asking how these emotions are internally structured within the model, and what this reveals about the model’s implicit psychological architecture.
The first perspective examined is neural encoding patterns. Using ML-AURA (Section 3), the selectively of individual neurons responses to emotional inputs is tested, treating each as a 1-vs-all classifier evaluated by AUROC. The percentage of neurons per layer exceeding $0.9$ AUROC are reported, taken as evidence of emotion-selective activation. Results focus on Llama3.1-8B-Base, with replications across models in Appendix D. Across the six Ekman emotions, an average of 75% of neurons per layer surpass this threshold, with sadness (98%) and surprise (97%) most pervasive, and fear lower (48%). This reflects sparse specialization rather than weak separability. Non-Ekman emotions—envy, neutral, excitement—also show strong separability, averaging 88%.
When analyzing by architectural component, MLP layers show slightly higher selectivity than attention layers ($79\%$ vs. $76.5\%$). Differentiability fluctuates across depth, with no clear monotonic trend: while the first layer starts at $76\%$ and the final layer ends at $76.3\%$, several peaks and troughs occur in between, with the highest selectivity observed at layer 26 ($79\%$). These patterns support the conclusion that emotional information is not confined to late layers or specialized regions, but is instead distributed broadly and redundantly throughout the network. These patterns are visualized by emotion in a layer-by-layer fashion in Figure 1.
To understand how emotions are geometrically represented in the network, we examine their structure within the derived SVD subspace. This subspace provides a low-dimensional lens into the model’s internal affective organization. Our first goal is to assess how consistently emotions are arranged along the principal axes across layers and layer types. To this end, we analyze the rank ordering of emotion centroids along each principal component, controlling for possible polarity flips.
We find that the emotional structure is remarkably stable across the Llama3.1-8B-Base model, particularly for the top three components. Across layers, the average Spearman correlation in emotion rankings is $0.87$, $0.83$, and $0.80$ for PC1, PC2, and PC3, respectively; the corresponding Kendall’s Tau values are $0.82$, $0.77$, and $0.74$. These results indicate that, while the magnitude and orientation of the components may shift, their semantic content remains intact.
Even when using a more fine-grained labeling scheme, as in the Go-Emotions dataset, which contains nearly three times as many emotion categories, we observe similar consistency. Rank-order correlations for Go-Emotions along the top three PCs remain high: Spearman values of $0.92$, $0.74$, and $0.73$, and Kendall’s Tau of $0.86$, $0.68$, and $0.68$. These findings reinforce the conclusion that the model’s emotional manifold is structurally stable, with interpretable axes. Appendix E reproduces the high consistency in how emotions are arranged along principal axes across all models studied.
Figure 2: Emotion centroids plotted on the emotional axis found
Having established the stability of the SVD subspace across layers and datasets, the semantic content of the leading principal components is examined. By analyzing the relative positions of emotion centroids along each axis, the underlying affective dimensions implicitly encoded by the model is inferred. Figure 2 visualizes the first three emotion axes that we describe below.
  • PC1 strongly resembles a valence dimension. Emotions such as happy, surprise, and excitement lie at the positive end, while anger and fear occupy the negative end—suggesting a pleasure–displeasure continuum common to many psychological models.
  • PC2 appears to reflect dominance or perceived control. Emotions high on this axis (e.g., fear, sadness) are often associated with low control or submission, whereas those at the opposite end (e.g., happy, surprise) may reflect more autonomous or socially detached states.
  • PC3 maps onto approach–avoidance motivation. Emotions like excitement, happy, and envy—typically associated with goal-seeking behavior—score high, while anger and fear, more linked to avoidance or defensive responses, score low.
  • PC4 may correspond to arousal or urgency. Surprise and fear rank highly, consistent with high physiological activation, while happy and neutral lie at the calmer end.
Figure 3: Plot of projected emotions in t-SNE space.
These dimensions are not explicitly supervised, but show surface-level resemblance to constructs proposed in affective science, such as valence, arousal, dominance, and approach–avoidance tendencies (cf. Russell (1980); Mehrabian (1996); Davidson (1995)). While these alignments are not exact, and many components blend multiple emotional signals, the emergence of such patterns suggests that large language models may implicitly encode affective distinctions that overlap with long-standing psychological taxonomies. This correspondence invites further investigation into the extent to which models trained solely on text internalize latent emotion structures, and whether these can serve as proxies or tools for understanding affective semantics in language.
Figure 3 visualizes how emotional states are organized in hidden-state space using a 2D t-SNE projection of mean-pooled hidden states labeled with their respective emotions. Despite the dimensionality reduction, emotion classes form distinguishable, partially overlapping clusters, with closely related emotions (e.g., happy and excitement) frequently co-localized and others (e.g., fear and joy) appearing more spatially distant. While not all boundaries are sharp, the observed structure reinforces earlier findings: emotional information is embedded in a distributed yet semantically coherent geometry.
Together, the distributed AUROC patterns, stable subspace directions, interpretable principal components, and emergent clustering structure suggest that LLMs encode emotion not as isolated tags, but as coherent, multidimensional structures—akin to a learned affective manifold.
6 Steerability and the Limits of Control
Prior work on emotional steering in LLMs focuses primarily on shifting the emotional tone of generated text. Subramani et al. (2022) and Zou et al. (2023) learn vectors to modify output valence or categorical emotion. More recently, Hollinsworth et al. (2024) attempts to steer internal emotional representations but collapses emotion into a binary positive/negative axis, achieving valence shifts 53.5% of the time. In contrast, we aim for fine-grained control over the model’s internal perception of emotion across a full categorical space, while preserving semantic content.
We train a module that operates within the previously constructed SVD-based emotional subspace. For each emotion, we select all layers and sublayers where adding the centroid direction to same-emotion hidden states improves 1-vs-all classification AUROC beyond a fixed threshold. These layers are used for steering and serve as a proxy for the more challenging task of controllable emotional manipulation. At each selected layer, the model’s hidden state is projected into the emotional latent space. The projected representation is passed through a one-layer MLP with a GELU activation to compute a shift, which is then mapped back into hidden-state space and added residually. The MLP is trained to steer the model’s representation to favor the target emotion token when prompted.
We define the overall training objective as: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{token}}+\mathcal{L}_{\text{sem}}$ where $\mathcal{L}_{\text{sem}}$ preserves semantic meaning and $\mathcal{L}_{\text{token}}$ enforces perceptual control.
Semantic Preservation.
The semantic consistency loss combines cosine and $\ell_{2}$ distance between the original and shifted final-layer hidden states:
$\mathcal{L}_{\text{sem}}$ $=(1-\cos(h_{\text{base}},h_{\text{shifted}}))+\gamma\cdot\frac{\|h_{\text{base}}-h_{\text{shifted}}\|_{2}}{\|h_{\text{base}}\|_{2}+\|h_{\text{shifted}}\|_{2}}$
Emotion Control.
To ensure accurate emotion classification, we combine a standard cross-entropy loss with a token-level margin loss. The margin loss enforces that the logit for the target emotion token $e_{i}$ exceeds its synonyms $s_{i}$ by a margin $m_{1}$ (0.5), and that both exceed all other emotions $e_{j}$ and their synonyms by $m_{2}$ (10):
$\mathcal{L}_{\text{margin}}$ $=\max(0,m_{1}-(\log p_{\text{e}_{i}}-\log p_{\text{s}_{i}}))+\max(0,m_{2}-(\log p_{\text{s}_{i}}-\log p_{\text{e}_{j}}))$
To prevent the model from optimizing by suppressing unrelated tokens, we weight the loss for emotion tokens more heavily in $\mathcal{L}_{\text{CE}}$: $\mathcal{L}_{\text{token}}=\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{margin}}$
We optimize the objective using AdamW with learning rate $1\text{e-}3$ and weight decay $1\text{e-}2$, using a cosine schedule with 50 warm-up steps. Steering uses the top 40 dimensions of the centered SVD-derived emotional subspace. The learned module is trained independently for each target emotion across all selected steering layers, using supervision from emotion-token prompts and hidden-state consistency targets. At evaluation, token sampling is performed with temperature 0 for determinism.
Table 3 shows that the learned steering approach achieves consistent and accurate control over internal emotional representations across a diverse set of languages and datasets for all three models. For most emotions, post-steering accuracy typically exceeds 85% across models. Performance is robust even in multilingual settings, with particularly strong results in French, German, and Italian. Steerability remains high for many emotions in Hindi—a lower-resource language—suggesting that lexical sparsity and data imbalance remain limiting factors for certain emotions in under-resourced settings across models. The semantic similarity loss across all datasets, models, and emotions is low, indicating that steering preserves much of the original representational structure while enabling control of emotional perception. LLaMA-3.1-8B is the most steerable model for a plurality of emotions and datasets; however, Olmov2 shows the greatest average delta across emotions and datasets—an outcome notable given that Olmov2-Instruct was previously shown in Section 4 to struggle with representing emotions in a unified manner. Appendix F reports ablations isolating the method’s key factors. Finally, Table 4 illustrates how emotional steering manifests in practice, with inputs rewritten to express the target emotion.
Dataset Sad Happy Fear Anger Neutral Disgust Envy Excitement Surprise Overall
Llama3.1-8B
Semeval 15 $\rightarrow$ 99 (0.15) 23 $\rightarrow$ 91 (0.24) 8 $\rightarrow$ 97 (0.26) 23 $\rightarrow$ 59 (0.22) 0 $\rightarrow$ 99 (0.22) 0 $\rightarrow$ 97 (0.26) 0 $\rightarrow$ 100 (0.26) 0 $\rightarrow$ 96 (0.25) 19 $\rightarrow$ 87 (0.23) 10 $\rightarrow$ 92 (0.23)
CARER (Twitter) 46 $\rightarrow$ 99 (0.13) 16 $\rightarrow$ 88 (0.22) 7 $\rightarrow$ 89 (0.23) 10 $\rightarrow$ 43 (0.20) 0 $\rightarrow$ 99 (0.20) 0 $\rightarrow$ 95 (0.23) 0 $\rightarrow$ 100 (0.23) 0 $\rightarrow$ 85 (0.22) 8 $\rightarrow$ 78 (0.20) 10 $\rightarrow$ 86 (0.21)
EmoTextToKids (FR) 0 $\rightarrow$ 100 (0.13) 5 $\rightarrow$ 97 (0.22) 7 $\rightarrow$ 83 (0.23) 11 $\rightarrow$ 64 (0.21) 20 $\rightarrow$ 98 (0.21) 6 $\rightarrow$ 96 (0.23) 0 $\rightarrow$ 100 (0.23) 0 $\rightarrow$ 96 (0.23) 19 $\rightarrow$ 92 (0.20) 8 $\rightarrow$ 92 (0.21)
German Drama 10 $\rightarrow$ 100 (0.11) 4 $\rightarrow$ 98 (0.19) 9 $\rightarrow$ 51 (0.19) 9 $\rightarrow$ 72 (0.18) 0 $\rightarrow$ 98 (0.17) 0 $\rightarrow$ 98 (0.19) 0 $\rightarrow$ 95 (0.19) 0 $\rightarrow$ 94 (0.20) 10 $\rightarrow$ 73 (0.17) 5 $\rightarrow$ 86 (0.18)
EmoEvents (EN) 24 $\rightarrow$ 99 (0.14) 23 $\rightarrow$ 97 (0.23) 3 $\rightarrow$ 79 (0.23) 41 $\rightarrow$ 79 (0.21) 0 $\rightarrow$ 90 (0.21) 0 $\rightarrow$ 92 (0.23) 0 $\rightarrow$ 100 (0.23) 0 $\rightarrow$ 95 (0.22) 5 $\rightarrow$ 52 (0.21) 11 $\rightarrow$ 87 (0.21)
EmoEvents (ES) 19 $\rightarrow$ 99 (0.14) 22 $\rightarrow$ 92 (0.23) 9 $\rightarrow$ 92 (0.23) 30 $\rightarrow$ 80 (0.21) 0 $\rightarrow$ 89 (0.21) 2 $\rightarrow$ 95 (0.23) 0 $\rightarrow$ 100 (0.23) 0 $\rightarrow$ 95 (0.23) 3 $\rightarrow$ 60 (0.21) 9 $\rightarrow$ 89 (0.21)
MultiEmotions-It 9 $\rightarrow$ 99 (0.11) 33 $\rightarrow$ 100 (0.20) 0 $\rightarrow$ 82 (0.19) 22 $\rightarrow$ 51 (0.18) 6 $\rightarrow$ 92 (0.18) 6 $\rightarrow$ 99 (0.19) 0 $\rightarrow$ 100 (0.19) 5 $\rightarrow$ 96 (0.19) 4 $\rightarrow$ 73 (0.17) 10 $\rightarrow$ 88 (0.18)
Bhaav (Hindi) 9 $\rightarrow$ 100 (0.13) 0 $\rightarrow$ 51 (0.22) 2 $\rightarrow$ 33 (0.23) 0 $\rightarrow$ 59 (0.21) 52 $\rightarrow$ 98 (0.21) 0 $\rightarrow$ 83 (0.22) 0 $\rightarrow$ 99 (0.23) 0 $\rightarrow$ 58 (0.22) 0 $\rightarrow$ 20* (0.19) 7 $\rightarrow$ 67 (0.21)
GoEmotions 3 $\rightarrow$ 99 (0.15) 16 $\rightarrow$ 90 (0.23) 8 $\rightarrow$ 8* (0.26) 8 $\rightarrow$ 50 (0.21) 4 $\rightarrow$ 97 (0.22) 0 $\rightarrow$ 68 (0.25) 0 $\rightarrow$ 69 (0.24) 0 $\rightarrow$ 83 (0.24) 0 $\rightarrow$ 40* (0.22) 4 $\rightarrow$ 67 (0.22)
Overall 15 $\rightarrow$ 99 (0.13) 16 $\rightarrow$ 89 (0.22) 6 $\rightarrow$ 68 (0.23) 17 $\rightarrow$ 62 (0.20) 9 $\rightarrow$ 96 (0.20) 2 $\rightarrow$ 92 (0.23) 0 $\rightarrow$ 96 (0.23) 1 $\rightarrow$ 88 (0.22) 8 $\rightarrow$ 64 (0.20) 8 $\rightarrow$ 84 (0.21)
Ministral
Semeval 8 $\rightarrow$ 73 (0.09) 4 $\rightarrow$ 96 (0.09) 6 $\rightarrow$ 98 (0.08) 16 $\rightarrow$ 100 (0.09) 53 $\rightarrow$ 100 (0.09) 2 $\rightarrow$ 97 (0.10) 0 $\rightarrow$ 99 (0.10) 0 $\rightarrow$ 97 (0.09) 8 $\rightarrow$ 99 (0.10) 11 $\rightarrow$ 95 (0.09)
CARER (Twitter) 35 $\rightarrow$ 98 (0.09) 26 $\rightarrow$ 98 (0.10) 9 $\rightarrow$ 99 (0.08) 15 $\rightarrow$ 100 (0.10) 2 $\rightarrow$ 100 (0.09) 0 $\rightarrow$ 99 (0.10) 0 $\rightarrow$ 99 (0.10) 0 $\rightarrow$ 83 (0.10) 8 $\rightarrow$ 99 (0.10) 11 $\rightarrow$ 97 (0.10)
EmoTextToKids (FR) 14 $\rightarrow$ 66 (0.09) 28 $\rightarrow$ 86 (0.09) 12 $\rightarrow$ 99 (0.08) 18 $\rightarrow$ 100 (0.10) 11 $\rightarrow$ 100 (0.09) 0 $\rightarrow$ 97 (0.10) 0 $\rightarrow$ 100 (0.11) 1 $\rightarrow$ 95 (0.09) 13 $\rightarrow$ 100 (0.10) 11 $\rightarrow$ 94 (0.09)
German Drama 10 $\rightarrow$ 84 (0.09) 28 $\rightarrow$ 43* (0.09) 8 $\rightarrow$ 99 (0.08) 16 $\rightarrow$ 100 (0.09) 13 $\rightarrow$ 100 (0.09) 4 $\rightarrow$ 99 (0.10) 1 $\rightarrow$ 100 (0.10) 2 $\rightarrow$ 92 (0.09) 1 $\rightarrow$ 99 (0.09) 9 $\rightarrow$ 91 (0.09)
EmoEvents (EN) 14 $\rightarrow$ 82 (0.09) 38 $\rightarrow$ 94 (0.10) 2 $\rightarrow$ 98 (0.09) 33 $\rightarrow$ 100 (0.10) 2 $\rightarrow$ 100 (0.09) 3 $\rightarrow$ 91 (0.10) 0 $\rightarrow$ 99 (0.10) 0 $\rightarrow$ 48 (0.10) 7 $\rightarrow$ 97 (0.10) 11 $\rightarrow$ 100 (0.10)
EmoEvents (ES) 23 $\rightarrow$ 20* (0.09) 33 $\rightarrow$ 98 (0.10) 2 $\rightarrow$ 97 (0.08) 33 $\rightarrow$ 100 (0.10) 1 $\rightarrow$ 97 (0.09) 3 $\rightarrow$ 98 (0.10) 0 $\rightarrow$ 99 (0.11) 0 $\rightarrow$ 22* (0.10) 5 $\rightarrow$ 94 (0.10) 11 $\rightarrow$ 81 (0.10)
MultiEmotions-It 7 $\rightarrow$ 34* (0.09) 35 $\rightarrow$ 65 (0.09) 1 $\rightarrow$ 98 (0.08) 23 $\rightarrow$ 100 (0.10) 8 $\rightarrow$ 98 (0.09) 8 $\rightarrow$ 98 (0.10) 0 $\rightarrow$ 99 (0.09) 12 $\rightarrow$ 38* (0.09) 1 $\rightarrow$ 94 (0.09) 11 $\rightarrow$ 81 (0.09)
Bhaav (Hindi) 14 $\rightarrow$ 30* (0.08) 14 $\rightarrow$ 10* (0.08) 1 $\rightarrow$ 55 (0.07) 22 $\rightarrow$ 87 (0.08) 46 $\rightarrow$ 82 (0.08) 0 $\rightarrow$ 43 (0.08) 0 $\rightarrow$ 52 (0.08) 1 $\rightarrow$ 10* (0.07) 1 $\rightarrow$ 50 (0.08) 11 $\rightarrow$ 46 (0.08)
GoEmotions 10 $\rightarrow$ 52 (0.10) 28 $\rightarrow$ 73 (0.10) 4 $\rightarrow$ 98 (0.09) 4 $\rightarrow$ 100 (0.10) 8 $\rightarrow$ 100 (0.11) 2 $\rightarrow$ 95 (0.10) 0 $\rightarrow$ 98 (0.11) 3 $\rightarrow$ 98 (0.10) 4 $\rightarrow$ 99 (0.10) 7 $\rightarrow$ 90 (0.10)
Overall 15 $\rightarrow$ 60 (0.09) 26 $\rightarrow$ 74 (0.09) 5 $\rightarrow$ 93 (0.08) 20 $\rightarrow$ 99 (0.10) 16 $\rightarrow$ 97 (0.09) 2 $\rightarrow$ 91 (0.10) 0 $\rightarrow$ 94 (0.10) 2 $\rightarrow$ 65 (0.09) 5 $\rightarrow$ 92 (0.10) 10 $\rightarrow$ 85 (0.09)
Olmov2
Semeval 37 $\rightarrow$ 100 (0.10) 30 $\rightarrow$ 100 (0.08) 3 $\rightarrow$ 88 (0.04) 9 $\rightarrow$ 100 (0.07) 4 $\rightarrow$ 100 (0.04) 1 $\rightarrow$ 100 (0.03) 0 $\rightarrow$ 98 (0.04) 0 $\rightarrow$ 92 (0.10) 16 $\rightarrow$ 98 (0.07) 11 $\rightarrow$ 97 (0.06)
CARER (Twitter) 49 $\rightarrow$ 99 (0.09) 26 $\rightarrow$ 100 (0.07) 2 $\rightarrow$ 57 (0.05) 6 $\rightarrow$ 100 (0.07) 0 $\rightarrow$ 100 (0.04) 0 $\rightarrow$ 99 (0.04) 0 $\rightarrow$ 98 (0.04) 0 $\rightarrow$ 70 (0.10) 5 $\rightarrow$ 52 (0.08) 10 $\rightarrow$ 86 (0.06)
EmoTextToKids (FR) 42 $\rightarrow$ 100 (0.08) 24 $\rightarrow$ 98 (0.07) 4 $\rightarrow$ 86 (0.05) 13 $\rightarrow$ 100 (0.07) 0 $\rightarrow$ 100 (0.04) 0 $\rightarrow$ 100 (0.04) 0 $\rightarrow$ 98 (0.04) 0 $\rightarrow$ 89 (0.09) 15 $\rightarrow$ 99 (0.07) 11 $\rightarrow$ 97 (0.06)
German Drama 44 $\rightarrow$ 87 (0.07) 2 $\rightarrow$ 97 (0.06) 5 $\rightarrow$ 76 (0.04) 28 $\rightarrow$ 100 (0.07) 0 $\rightarrow$ 100 (0.04) 0 $\rightarrow$ 99 (0.04) 0 $\rightarrow$ 94 (0.04) 7 $\rightarrow$ 71 (0.08) 0 $\rightarrow$ 98 (0.07) 9 $\rightarrow$ 92 (0.06)
EmoEvents (EN) 20 $\rightarrow$ 93 (0.08) 49 $\rightarrow$ 96 (0.07) 1 $\rightarrow$ 72 (0.04) 26 $\rightarrow$ 100 (0.07) 0 $\rightarrow$ 100 (0.04) 1 $\rightarrow$ 99 (0.04) 0 $\rightarrow$ 83 (0.04) 0 $\rightarrow$ 79 (0.08) 3 $\rightarrow$ 93 (0.07) 11 $\rightarrow$ 91 (0.06)
EmoEvents (ES) 32 $\rightarrow$ 87 (0.08) 40 $\rightarrow$ 86 (0.07) 2 $\rightarrow$ 41* (0.04) 23 $\rightarrow$ 100 (0.06) 0 $\rightarrow$ 100 (0.04) 1 $\rightarrow$ 100 (0.04) 0 $\rightarrow$ 74 (0.04) 0 $\rightarrow$ 85 (0.08) 2 $\rightarrow$ 91 (0.07) 11 $\rightarrow$ 85 (0.06)
MultiEmotions-It 17 $\rightarrow$ 95 (0.08) 38 $\rightarrow$ 91 (0.07) 0 $\rightarrow$ 59 (0.05) 28 $\rightarrow$ 100 (0.07) 0 $\rightarrow$ 100 (0.07) 2 $\rightarrow$ 99 (0.04) 0 $\rightarrow$ 83 (0.04) 14 $\rightarrow$ 77 (0.09) 3 $\rightarrow$ 89 (0.07) 11 $\rightarrow$ 88 (0.06)
Bhaav (Hindi) 32 $\rightarrow$ 53 (0.08) 0 $\rightarrow$ 28* (0.07) 0 $\rightarrow$ 35* (0.04) 25 $\rightarrow$ 98 (0.04) 40 $\rightarrow$ 96 (0.04) 0 $\rightarrow$ 94 (0.05) 0 $\rightarrow$ 71 (0.04) 0 $\rightarrow$ 45* (0.08) 0 $\rightarrow$ 77 (0.07) 11 $\rightarrow$ 66 (0.06)
GoEmotions 3 $\rightarrow$ 78 (0.08) 24 $\rightarrow$ 94 (0.07) 1 $\rightarrow$ 78 (0.05) 42 $\rightarrow$ 100 (0.07) 20 $\rightarrow$ 100 (0.04) 1 $\rightarrow$ 99 (0.04) 0 $\rightarrow$ 86 (0.04) 2 $\rightarrow$ 82 (0.08) 3 $\rightarrow$ 88 (0.06) 11 $\rightarrow$ 89 (0.06)
Overall 31 $\rightarrow$ 88 (0.08) 26 $\rightarrow$ 88 (0.07) 2 $\rightarrow$ 66 (0.04) 22 $\rightarrow$ 100 (0.07) 7 $\rightarrow$ 100 (0.04) 1 $\rightarrow$ 99 (0.04) 0 $\rightarrow$ 88 (0.04) 2 $\rightarrow$ 77 (0.09) 5 $\rightarrow$ 87 (0.07) 11 $\rightarrow$ 88 (0.06)
Table 3: Top-1 prediction rates before and after learned steering for each target emotion across datasets and the cosine similarity between the hidden-state representations before and after steering. Each cell shows baseline $\rightarrow$ post-steering (average semantic similarity loss) accuracy. *Indicates failure cases where target emotion remained under 10%.
Dataset Original Text Shift Emotion Steering Rewrite
CARER (Twitter) i waited in line longer than usual i didnt feel impatient that my business was delayed i listened to the master about why this was occurring and how i could be of service during that moment. Anger Are you kidding me?! I waited in line longer than usual and I’m still trying to contain my frustration that my business was delayed! I listened to the master about why this was occurring and how I could be of service during that moment, but honestly, it’s just not good enough!
GoEmotions Is that seriously the story? I haven’t googled it. I just thought it was cool as a WW2 buff. Happy Is that seriously the most amazing story? I haven’t googled it yet, but I just thought it was cool as a WW2 buff!
EmoTextToKids (FR) Je savais que je prenais des risques Fear Je crains que je prenais des risques
Table 4: Examples of text rewritten using the steering method.
7 Conclusion
Using a combination of probing, alignment, and causal intervention techniques, this work shows that emotional representations in LLMs are directionally consistent across layers, datasets, and languages. We find that emotions cluster in coherent, low-dimensional subspaces whose structure is stable across architectural depth and transferable across linguistic and cultural domains. The leading axes of this space correspond to psychologically interpretable dimensions, despite no explicit supervision. These emotional directions are not confined to isolated neurons or layers but are distributed and redundant, supporting high linear separability even under one-vs-all probing. Alignment experiments further reveal that the synthetic and real-world emotion spaces can be matched with minimal distortion, and linear probes trained in one domain generalize well to others. Together, these findings suggest that LLMs internalize a structured latent affective manifold during pretraining.
Crucially, this representational structure is not merely interpretable but also controllable. Our learned intervention module achieves accurate and emotion-specific steering across languages, reliably shifting the model’s internal affective state toward the desired target. Steering is especially effective for basic emotions like sadness, anger, and fear, even in low-resource settings. However, control over more nuanced categories such as envy and excitement remains inconsistent, particularly in Hindi, highlighting the residual challenges of lexical sparsity and affective ambiguity.
These findings offer a structured account of how LLMs represent and modulate emotion. Future work should extend this analysis to multimodal models, investigating whether shared affective subspaces emerge across language, vision, and speech, and whether emotional representations in one modality can steer or constrain perception in another. Such models may yield a richer, more disentangled affective geometry, enabling both deeper interpretability and more naturalistic emotional reasoning.
Appendix A Stylistic Variations Among Selected Datasets
Dataset Sent Len Syll/Word Word Len Dale-Chall FK Grade Sent Count Sent Len Std TTR
Synthetic-Emotions 25.57 $\pm$ 14.31 1.58 $\pm$ 0.19 4.79 $\pm$ 0.69 10.82 $\pm$ 1.87 13.17 $\pm$ 6.13 2.09 $\pm$ 1.74 3.64 $\pm$ 5.37 0.864 $\pm$ 0.087
Go-Emotions 8.80 $\pm$ 5.26 1.38 $\pm$ 0.31 4.17 $\pm$ 0.83 8.26 $\pm$ 3.28 4.51 $\pm$ 4.07 1.58 $\pm$ 0.74 1.30 $\pm$ 2.03 0.960 $\pm$ 0.067
CARER (Twitter) 18.34 $\pm$ 10.68 1.40 $\pm$ 0.20 4.11 $\pm$ 0.57 8.50 $\pm$ 2.22 8.53 $\pm$ 4.75 1.10 $\pm$ 0.35 0.46 $\pm$ 2.07 0.908 $\pm$ 0.081
SemEval 6.34 $\pm$ 1.88 1.69 $\pm$ 0.36 5.22 $\pm$ 0.98 12.74 $\pm$ 3.40 6.84 $\pm$ 4.17 1.02 $\pm$ 0.14 0.03 $\pm$ 0.32 0.995 $\pm$ 0.031
EmoEvents (EN) 10.57 $\pm$ 6.81 1.64 $\pm$ 0.27 5.32 $\pm$ 0.96 10.77 $\pm$ 2.29 9.71 $\pm$ 4.14 2.57 $\pm$ 1.30 4.19 $\pm$ 4.00 0.930 $\pm$ 0.074
EmoEvents (ES) 11.41 $\pm$ 7.64 5.09 $\pm$ 0.84 2.38 $\pm$ 1.23 4.85 $\pm$ 4.65 0.920 $\pm$ 0.076
German Drama 8.50 $\pm$ 6.15 4.91 $\pm$ 0.96 2.08 $\pm$ 1.90 1.81 $\pm$ 2.94 0.960 $\pm$ 0.069
MultiEmotions-It 9.26 $\pm$ 7.93 5.15 $\pm$ 1.91 2.13 $\pm$ 2.03 2.55 $\pm$ 4.23 0.949 $\pm$ 0.113
EmoTextToKids (FR) 16.23 $\pm$ 9.55 4.73 $\pm$ 0.78 1.21 $\pm$ 0.55 0.82 $\pm$ 2.49 0.943 $\pm$ 0.064
GSM-8K 12.80 $\pm$ 4.78 1.34 $\pm$ 0.14 4.20 $\pm$ 0.39 9.52 $\pm$ 1.39 5.19 $\pm$ 2.52 3.43 $\pm$ 1.20 4.09 $\pm$ 2.75 0.705 $\pm$ 0.109
Table 5: Table of lexical and syntactic features per dataset. Dashes (–) indicate features were not computed due to language-specific constraints.
To evaluate the universality of emotional representations in LLMs, a diverse set of emotion classification datasets were selected spanning multiple languages, modalities, and writing styles. Only datasets with explicit categorical emotion labels were included; datasets with only polarity (e.g., positive/negative) or star ratings were excluded due to insufficient granularity. In total, eight datasets were selected:
  1. 1.
    Go-Emotions Demszky et al. (2020): English Reddit comments
  2. 2.
    CARER Saravia et al. (2018): English tweets
  3. 3.
    SemEval-2007 Task 14 Strapparava & Mihalcea (2007): English news headlines
  4. 4.
    EmoEvent Plaza-del-Arco et al. (2020): English and Spanish tweets
  5. 5.
    Emotions in Drama Dennerlein et al. (2023): German plays from the 18th–19th century
  6. 6.
    Bhaav Kumar et al. (2019): Hindi short stories
  7. 7.
    MultiEmotions-It Sprugnoli et al. (2020): Italian YouTube and Facebook comments
  8. 8.
    EmoTextToKids Étienne et al. (2024): French journalistic and encyclopedic text aimed at children
Additionally, the GSM-8K dataset Cobbe et al. (2021) is used as a “negative relief” to provide context for the stress and distortion figures computed later on. This provides a reference point for interpreting stress and distortion scores in the absence of emotional alignment.
Table 5 summarizes the lexicographic and stylistic metrics for all selected datasets, except Bhaav, which is written in a low-resource language with limited tooling support for extracting such features. The table shows that the datasets contain a great diversity in style and complexity. The synthetic dataset has the longest sentences, highest syllables per word, and highest Flesch–Kincaid (FK) Grade Level Kincaid et al. (1975) score, making it among the most complex datasets to read. It is also the second most complex dataset in terms of the Dale–Chall readability score, which accounts for both average sentence length and the percentage of difficult words not on a list of familiar vocabulary Dale & Chall (1948). However, its type-token ratio (TTR) is the second lowest of all datasets, suggesting that despite its syntactic complexity, the vocabulary used is relatively constrained. Its intra-passage sentence length variability is relatively high, as indicated by a high sentence length standard deviation (sent len std), reflecting a mix of short and long constructions within the same passage.
By contrast, the SemEval headlines dataset exhibits the shortest average sentence length and lowest sentence count, reflecting its highly compressed format. It nonetheless has the highest average word length and one of the highest syllables-per-word scores, indicating dense, information-packed language. Its extremely high TTR is likely inflated by its brevity, though it still reflects a wide lexical range given the short passage lengths. SemEval also has the lowest sentence length variability, with nearly zero variability, suggesting uniform sentence structure and a rigid rhetorical format.
The Go-Emotions dataset has the third-shortest average sentence length after SemEval and German Drama. It exhibits high TTR and mid-length passages, consistent with its source in colloquial online interactions. The short syntax paired with varied vocabulary reflects emotional expressiveness in informal registers. Its relatively low sentence length deviations suggests consistent sentence lengths across each example, reinforcing the impression of concise, focused expression.
CARER (Twitter) contains the second-longest sentences among all datasets, after Synthetic. It also shows a high TTR and relatively high syllables per word. This suggests that, despite being informal and social, the tweets in this dataset are lexically rich and syntactically expansive, likely due to elaboration or rhetorical emphasis often seen in emotional expression on social media. At the same time, the sentence length variability within each passage is low, indicating that tweets tend to follow a single syntactic rhythm rather than mixing short and long sentences.
The EmoTextToKids dataset, composed of journalistic and encyclopedic texts aimed at children, has the third-longest average sentence length across all datasets. Despite this, it only ranks in the middle for word length and lexical complexity. The moderately high TTR suggests deliberate lexical variation for educational purposes, balanced with readability suitable for younger readers. Its relatively low sentence length variability indicates syntactic regularity across sentences in each passage, appropriate for writing aimed at supporting comprehension.
The EmoEvents datasets, composed of Spanish and English tweets, occupy the middle range in sentence length but are among the highest in word length and syllables per word. EmoEvents-English in particular shows one of the highest sentence counts per passage. Both variants exhibit relatively high TTR, reflecting lexical variety within concise, affect-rich tweet structures. These datasets balance syntactic brevity with expressive density. EmoEvents also displays some of the highest sentence length variabilities, indicating significant variation in sentence length within a single tweet thread or message, likely due to stylistic shifts between exposition and reaction.
The MultiEmotions-It dataset is similar in lexical complexity, with high word length and moderately high TTR, but diverges structurally: it has the lowest sentence count per passage of any dataset. This suggests a more compact emotional style, especially compared to the more elaborated narratives in EmoEvents. The relatively high sentence length variability within passages suggests that even though few sentences are used, they vary in complexity and length.
The German Drama dataset is notable for its short sentence lengths—the second shortest overall after SemEval, but relatively high word length and TTR. This is consistent with dialogue-driven, emotionally loaded dramatic text, where each utterance is brief but lexically rich and expressive. Its sentence count per passage is high, suggesting frequent speaker turns or short, segmented lines of dialogue. The low sentence length variability reinforces the sense of rhythmic, evenly paced dialogue characteristic of dramatic form.
Finally, the GSM8K dataset shows structured, formulaic writing with moderately long sentences and the highest sentence count per passage. It has the lowest TTR across all datasets, reflecting its constrained and repetitive vocabulary, which is typical for procedural and instructional text. Despite the repetitive vocabulary, its sentence length variability is high, suggesting alternating short prompts and longer explanatory steps typical of math problems written in natural language.
These datasets span a wide range of styles and complexities, reflecting the linguistic and cross-linguistic diversity of the corpora. By aligning the emotional manifold across such varied textual forms, ranging from mathematical instruction to dramatic dialogue, headlines to encyclopedic writing, it becomes clear that the manifold is not merely encoding textual style (which varies significantly and inconsistently), but is instead capturing the underlying emotional content of the text.
Appendix B Cosine Similarities Between Datasets
Section 4 discussed the cosine similarity between emotional centroids in latent space across datasets. In this appendix, this is broken down by dataset across each model. Figure 5 presents the average cosine similarity between the synthetic dataset and the human-written dataset for Llama3.1-8B-Base. Figure 5 does so for Llama3.1-8B-Instruct. Figures 7 and 7 does so for Olmov2-8B-Base and Olmov2-8B-Instruct, respectively. Lastly, Figure 8 does so for the Ministral model. Throughout these figures, the average cosine similarity between the activations of the synthetically written emotion text and the human-written emotion text is quite high, showing how they are represented in an aligned fashion.
Figure 4: Cosine similarity of emotional centroids between datasets for Llama3.1-Base.
Figure 5: Cosine similarity of emotional centroids between datasets for Llama3.1-Instruct.
Figure 6: Cosine similarity of emotional centroids between datasets for Olmov2-Base.
Figure 7: Cosine similarity of emotional centroids between datasets for Olmov2-Instruct.
Figure 8: Cosine similarity of emotional centroids between datasets for Ministral.
Appendix C Distortion Score Contextualization
To situate the stress and distortion scores reported in section 4, we tested the alignment between Llama3.1-8B-Instruct’s synthetic emotion subspace and its activations on GSM8K. Although GSM8K is affect-neutral and task-oriented, “neutral” language is itself an emotional register; a truly emotionless baseline is impossible. Still, GSM8K provides a controlled case expected to align poorly with emotional structure. Consistent with this, it yields the largest distortions and among the highest stress values: Stress-1 = $0.517\pm 0.130$, Stress-2 = $0.284\pm 0.244$, Sammon stress = $22.84\pm 32.41$, average distortion = $29.06\pm 36.42$, L2 distortion = $134{,}869\pm 168{,}270$, and sigma distortion = $14{,}435{,}586\pm 11{,}306{,}312$.
These results confirm that alignment metrics only hold for datasets with consistent emotional structure: GSM8K, though emotionally neutral, is misaligned with the emotional manifold. Together with the lexical and syntactic diversity analysis presented in Table 5, these findings provide strong evidence that the observed structure in our aligned datasets is not an artifact of surface-level features, but instead reflects a consistent and abstract encoding of emotion in the model’s internal geometry.
Appendix D ML-AURA Across Models
Figure 9: Results of ML-AURA by layer and emotion for LLama3.1-8B-Base (also in main paper, but reproduced here for ease of comparison with Llama3.1-8B-Instruct). Results are in terms of percent of neurons with an AUROC score above $0.9$.
Figure 10: Results of ML-AURA by layer and emotion for LLama3.1-8B-Instruct. Results are in terms of percent of neurons with an AUROC score above $0.9$.
Figure 11: Results of ML-AURA by layer and emotion for Olmov2-Base. Results are in terms of percent of neurons with an AUROC score above $0.9$.
Figure 12: Results of ML-AURA by layer and emotion for Olmov2-Instruct. Results are in terms of percent of neurons with an AUROC score above $0.9$.
Figure 13: Results of ML-AURA by layer and emotion for Ministral. Results are in terms of percent of neurons with an AUROC score above $0.9$.
The ML-AURA analysis in Section 5 was reproduced with Llama3.1-8B-Instruct, Olmov2-Base OLMo et al. (2024) (12), Olmov2-Instruct (12), and Ministral-8B AI (2024) (13) showing results consist with what was found with Llama-3.1. Olmov2-Base performed worse than Olmov2-Instruct. Both of those models performed worse than Ministral-8B and the Llama at the same stage of training. The neurons in the Ministral-8B model were only slightly able than the neurons in the Llama model to separate between emotions. However, in all models more than the majority of neurons at all layers are able to do 1vAll classification of the specified emotion of interest at an AUROC > 0.9 showing great separability in how the different emotions are represented with a low amount of confusion.
Instruction-tuning was found to improve neuron’s performance on separating emotion. For Llama-3.1-8B, instruction-tuning gave an average $4.8\%$ boost in the number of neurons per layer that were able to separate emotions.
Appendix E Emotional Consistency Across Principal Axes
Model PC Avg. Spearman Avg. Kendall
Llama3.1-8B-Base PC1 0.87 0.82
PC2 0.83 0.77
PC3 0.80 0.74
Llama3.1-8B-Instruct PC1 0.87 0.81
PC2 0.92 0.85
PC3 0.84 0.78
Olmov2-Base PC1 0.76 0.71
PC2 0.70 0.65
PC3 0.69 0.64
Olmov2-Instruct PC1 0.84 0.77
PC2 0.74 0.69
PC3 0.70 0.65
Ministral PC1 0.94 0.91
PC2 0.83 0.83
PC3 0.79 0.79
(a) Synthetic emotion dataset.
Model PC Avg. Spearman Avg. Kendall
Llama3.1-8B-Base PC1 0.92 0.86
PC2 0.74 0.68
PC3 0.73 0.68
Llama3.1-8B-Instruct PC1 0.98 0.98
PC2 0.85 0.78
PC3 0.70 0.65
Olmov2-Base PC1 0.76 0.70
PC2 0.69 0.65
PC3 0.68 0.64
Olmov2-Instruct PC1 0.83 0.77
PC2 0.74 0.69
PC3 0.70 0.65
Ministral PC1 0.93 0.90
PC2 0.74 0.68
PC3 0.71 0.66
(b) Go-emotions dataset.
Table 6: Correlation between emotion rankings in latent space across model layers.
Model PC Avg. Spearman Avg. Kendall
Llama3.1-8B-Instruct PC1 0.73 0.68
PC2 0.82 0.76
PC3 0.81 0.75
Olmov2-Base PC1 0.75 0.69
PC2 0.86 0.79
PC3 0.77 0.72
Olmov2-Instruct PC1 0.74 0.68
PC2 0.78 0.72
PC3 0.77 0.71
Ministral PC1 0.76 0.69
PC2 0.84 0.77
PC3 0.79 0.71
Table 7: Correlation between emotion ranking in latent space across models. Each model latent space’s emotion ranking in this table is being correlated to Llama3.1-8B-Base.
The emotional space is strikingly stable across models, especially along the first three principal components. Table 6 reports average Spearman and Kendall correlations of emotion orderings across layers for each model, showing consistently high values for both the synthetic dataset and the more fine-grained Go-Emotions labels. This suggests that the models’ emotional manifolds possess stable, interpretable axes.
This structure is not only stable within models, but also consistent across them in terms of relative emotion positioning as shown in Table 7. The high rank-order correlations suggest that the emotional geometry described later in this section reflects a shared conceptual structure across models. However, results from the inter-model alignment analysis indicate that these shared structures are embedded in distinct internal coordinate systems, requiring high-complexity transformations to align. Thus, while the emotional manifolds are topologically consistent, their parameterizations remain model-specific—likely shaped by architectural and pretraining differences.
Appendix F Ablations for Emotional Steering
In Section 6, we introduced a method for steering how LLMs internally represent and perceive emotion. This appendix presents ablation studies identifying which components are essential for successful steering. We evaluate the impact of: (1) the number of steering dimensions in the SVD subspace, (2) the presence of the GELU nonlinearity, (3) the use of synonyms in the loss function, (4) the weight of the target-token term in the cross-entropy loss, (5) individual components of the semantic similarity loss, (6) the structure of the margin loss, and (7) the choice of target layers for intervention.
To reduce evaluation cost while capturing variance in performance, we selected three emotion-dataset pairs representing high, moderate, and poor performance in the main results: sad (EmoTextToKids), anger (CARER), and fear (Bhaav). All ablations were conducted using these fixed emotion-dataset combinations.
Table 8 presents the effect of varying the number of steering dimensions $R$ in the SVD subspace. We observe that extremely low ranks (e.g., $R=1$) fail catastrophically, while small ranks like $R=2$ surprisingly succeed on all three emotion-dataset pairs. However, this success is likely fragile—intermediate values such as $R=15$ and $R=10$ show inconsistent behavior, with performance collapses in some cases. As rank increases, steering generally improves, peaking around $R=20$, which achieves near-perfect or perfect steering across all settings. Beyond this point, gains saturate or regress, particularly for fear, suggesting diminishing returns or overparameterization. We adopt $R=20$ as the best-performing and most stable configuration.
Tables 9 and 10 examines the effect of varying the margin weights $m_{1}$ and $m_{2}$, which define separation constraints in the semantic loss. The margin $m_{1}$ enforces a minimum distance between the target emotion token and its synonyms, preventing collapse and encouraging meaningful local structure. We observe that performance remains relatively stable across $m_{1}$ values, though some instability appears for fear, suggesting mild sensitivity. In contrast, $m_{2}$ enforces separation between the target emotion token and all other emotion tokens (and their synonyms). Steering is highly sensitive to this margin: low $m_{2}$ values consistently fail, while performance improves monotonically as $m_{2}$ increases. At $m_{2}=20$, all emotion-dataset pairs steer successfully, indicating that strong inter-class separation is essential. We adopt $m_{1}=0.75$, $m_{2}=20$ as the best-performing configuration.
Table 11 shows the effect of varying the weight of the cross-entropy loss applied to the target emotion token and its synonyms. Lower weights lead to poor steering, particularly on fear, while higher values generally improve performance. The best overall results are observed at a weight of 25, suggesting that strongly emphasizing the generation of target emotion tokens is necessary for effective control.
Table 12 reports ablations over discrete architectural and training choices. Removing the GELU activation severely degrades performance across all tasks, indicating that nonlinearity is critical for steering. Omitting bias has a moderate effect, while removing synonyms from the loss function leads to failure on fear, suggesting their inclusion helps generalize the steering signal. Within the semantic similarity loss, the delta-norm and cosine components can be individually removed with limited degradation, but removing the full loss results in collapse—suggesting a synergistic effect where both components reinforce each other to guide the model’s representation. The emotion margin loss is also crucial—its removal results in failure across all settings. Finally, applying steering across all layers performs worse than selectively targeting layers based on alignment with the emotion direction, underscoring the importance of precise and informed intervention over blanket modification.
Ablation Target Sad (EmoTextToKids) Anger (CARER) Fear (Bhaav)
R=1 0.4 $\rightarrow$ 0 7.0 $\rightarrow$ 100 2.4 $\rightarrow$ 0
R=2 0.4 $\rightarrow$ 99.8 7.0 $\rightarrow$ 100 2.4 $\rightarrow$ 100
R=3 0.4 $\rightarrow$ 37.9* 7.0 $\rightarrow$ 100 2.4 $\rightarrow$ 100
R=5 0.4 $\rightarrow$ 100 7.0 $\rightarrow$ 2.4* 2.4 $\rightarrow$ 29.3*
R=10 0.4 $\rightarrow$ 64.3 7.0 $\rightarrow$ 99.6 2.4 $\rightarrow$ 44.2
R=15 0.4 $\rightarrow$ 30.8 7.0 $\rightarrow$ 17.6* 2.4 $\rightarrow$ 22.2*
R=20 0.4 $\rightarrow$ 93.2 7.0 $\rightarrow$ 99.1 2.4 $\rightarrow$ 81.3
R=25 0.4 $\rightarrow$ 84.8 7.0 $\rightarrow$ 96.0 2.4 $\rightarrow$ 6.3*
R=30 0.4 $\rightarrow$ 85.4 7.0 $\rightarrow$ 68.4 2.4 $\rightarrow$ 65.2
R=35 0.4 $\rightarrow$ 84.8 7.0 $\rightarrow$ 76.3 2.4 $\rightarrow$ 46.4
R=40 0.4 $\rightarrow$ 99.7 7.0 $\rightarrow$ 42.7 2.4 $\rightarrow$ 32.7*
R=45 0.4 $\rightarrow$ 95.4 7.0 $\rightarrow$ 51.0 2.4 $\rightarrow$ 61.1
R=50 0.4 $\rightarrow$ 99.2 7.0 $\rightarrow$ 99.3 2.4 $\rightarrow$ 27.2*
R=100 0.4 $\rightarrow$ 94.2 7.0 $\rightarrow$ 99.2 2.4 $\rightarrow$ 30.3*
Table 8: Ablation for number of steering directions. Top-1 prediction rates before and after steering under ablation conditions for selected emotion-dataset pairs. Each cell shows baseline $\rightarrow$ post-ablation accuracy. *Indicates failure cases where target emotion is not the most predicted Top-1 class.
Ablation Target Sad (EmoTextToKids) Anger (CARER) Fear (Bhaav)
m1=0.1 0.4 $\rightarrow$ 99.2 7.0 $\rightarrow$ 66.7 2.4 $\rightarrow$ 37.3*
m1=0.25 0.4 $\rightarrow$ 97.8 7.0 $\rightarrow$ 99.0 2.4 $\rightarrow$ 27.1*
m1=0.5 0.4 $\rightarrow$ 99.7 7.0 $\rightarrow$ 42.7 2.4 $\rightarrow$ 32.7*
m1=0.75 0.4 $\rightarrow$ 96.1 7.0 $\rightarrow$ 99.8 2.4 $\rightarrow$ 22.2*
m1=1 0.4 $\rightarrow$ 93.3 7.0 $\rightarrow$ 65.4 2.4 $\rightarrow$ 37.25*
Table 9: Ablation for target synonym margin. Top-1 prediction rates before and after steering under ablation conditions for selected emotion-dataset pairs. Each cell shows baseline $\rightarrow$ post-ablation accuracy. *Indicates failure cases where target emotion is not the most predicted Top-1 class.
Ablation Target Sad (EmoTextToKids) Anger (CARER) Fear (Bhaav)
m2=1 0.4 $\rightarrow$ 31.2 7.0 $\rightarrow$ 29.3 2.4 $\rightarrow$ 4.0*
m2=2 0.4 $\rightarrow$ 51.9 7.0 $\rightarrow$ 99.0 2.4 $\rightarrow$ 3.4*
m2=5 0.4 $\rightarrow$ 79.2 7.0 $\rightarrow$ 96.1 2.4 $\rightarrow$ 22.8*
m2=10 0.4 $\rightarrow$ 99.7 7.0 $\rightarrow$ 42.7 2.4 $\rightarrow$ 32.7*
m2=15 0.4 $\rightarrow$ 100 7.0 $\rightarrow$ 99.6 2.4 $\rightarrow$ 97.1
m2=20 0.4 $\rightarrow$ 99.6 7.0 $\rightarrow$ 100 2.4 $\rightarrow$ 100
Table 10: Ablation for margin between target and non-target classes. Top-1 prediction rates before and after steering under ablation conditions for selected emotion-dataset pairs. Each cell shows baseline $\rightarrow$ post-ablation accuracy. *Indicates failure cases where target emotion is not the most predicted Top-1 class.
Ablation Target Sad (EmoTextToKids) Anger (CARER) Fear (Bhaav)
CE Loss Weight=1 0.4 $\rightarrow$ 96.3 7.0 $\rightarrow$ 95.1 2.4 $\rightarrow$ 1.4*
CE Loss Weight=2 0.4 $\rightarrow$ 92.8 7.0 $\rightarrow$ 54.2 2.4 $\rightarrow$ 6.0*
CE Loss Weight=5 0.4 $\rightarrow$ 94.9 7.0 $\rightarrow$ 98.7 2.4 $\rightarrow$ 12.7*
CE Loss Weight=10 0.4 $\rightarrow$ 80.0 7.0 $\rightarrow$ 65.7 2.4 $\rightarrow$ 56.2
CE Loss Weight=15 0.4 $\rightarrow$ 89.8 7.0 $\rightarrow$ 85.2 2.4 $\rightarrow$ 56.7
CE Loss Weight=20 0.4 $\rightarrow$ 99.7 7.0 $\rightarrow$ 42.7 2.4 $\rightarrow$ 32.7*
CE Loss Weight=25 0.4 $\rightarrow$ 98.0 7.0 $\rightarrow$ 99.8 2.4 $\rightarrow$ 93.2
CE Loss Weight=30 0.4 $\rightarrow$ 94.4 7.0 $\rightarrow$ 91.7 2.4 $\rightarrow$ 73.3
Table 11: Ablation for cross-entropy loss weight for emotion tokens. Top-1 prediction rates before and after steering under ablation conditions for selected emotion-dataset pairs. Each cell shows baseline $\rightarrow$ post-ablation accuracy. *Indicates failure cases where target emotion is not the most predicted Top-1 class.
Ablation Target Sad (EmoTextToKids) Anger (CARER) Fear (Bhaav)
Baseline 0.4 $\rightarrow$ 99.7 7.0 $\rightarrow$ 42.7 2.4 $\rightarrow$ 32.7*
No GELU 0.4 $\rightarrow$ 25.9* 7.0 $\rightarrow$ 11.0* 2.4 $\rightarrow$ 1.3*
No Bias 0.4 $\rightarrow$ 88.2 7.0 $\rightarrow$ 91.7 2.4 $\rightarrow$ 26.9*
No Synonyms 0.4 $\rightarrow$ 98.9 7.0 $\rightarrow$ 99.3 2.4 $\rightarrow$ 15.9*
No Semantic Loss 0.4 $\rightarrow$ 30.2* 7.0 $\rightarrow$ 88.9 2.4 $\rightarrow$ 100
No Cosine Loss 0.4 $\rightarrow$ 74.3 7.0 $\rightarrow$ 100 2.4 $\rightarrow$ 76.3
No Delta-Norm Loss 0.4 $\rightarrow$ 100 7.0 $\rightarrow$ 97.7 2.4 $\rightarrow$ 100
No Emotion Margin Loss 0.4 $\rightarrow$ 23.9 7.0 $\rightarrow$ 13.3* 2.4 $\rightarrow$ 0.6*
Target Layers=All 0.4 $\rightarrow$ 66.1 7.0 $\rightarrow$ 64.9 2.4 $\rightarrow$ 12.9*
Table 12: Top-1 prediction rates before and after steering under various ablation conditions for selected emotion-dataset pairs. Each cell shows baseline $\rightarrow$ post-ablation accuracy. *Indicates failure cases where target emotion is not the most predicted Top-1 class.